OpenStudy (anonymous):

PC X takes 0:13:15 (h:mm:ss) to render a 5 minute video PC Y takes 0:04:31 (h:mm:ss) to render the same 5 minute video How much faster is PC Y?

OpenStudy (anonymous):

do the calculation in seconds. (13x60)+15=795=PC x PC Y = (4x60)+31 = 271s is the answer in percentage ? 271/795 x 100 = 34.1% PC y is 34 % faster

OpenStudy (anonymous):

First, Commit showed that PC Y takes 34.1% as long as PC X to do something. That doesn't mean it's 34.1% faster. You could say it's 34.1% as slow as PC X, but that sounds weird. I don't like the wording of the question because it doesn't say what it means by "how much faster." We could give an absolute difference in the amount of work done per second or minute. Or we could give a relative multiplier of their speeds and say PC Y is n times as fast as PC X. Or say how much PC Y's speed exceeds PC X's speed, in relation to PC X's speed. 1. PC X takes 795s to do 1 unit of work PC Y takes 271s to do 1 unit of work PC X's speed = Sx = 1 unit / 795s = 0.001258 u/s PC Y's speed = Sy = 1 unit / 271s = 0.00369 u/s So, PC Y is 0.002432u/s faster than PC X. 2. Sy/Sx = (.00369u/s) / (.001258u/s) = 2.933 = 293.3% So PC Y is 2.933 times as fast as PC X. Or PC Y is 293.3% as fast as PC X. 3. (Sy - Sx)/Sx = Sy/Sx - 1 = 2.933-1 = 1.933 = 193.3% So PC Y is 1.933 times faster than PC X. Or PC Y is 193.3% faster than PC X. I'm guessing #3 is probably the most appropriate answer.